# OpenECAD: An Efficient Visual Language Model for Computer-Aided Design

### v2

@@ 6.5.1 Refinement
At the same time, we have added a Constraints section to the dataset. We consider nine types of widely used constraint commands (see Table 5): making a line horizontal or vertical, fixing the size of a curve, making two points of curves coincident, making two lines parallel, perpendicular, or at a fixed angle, and making two curves tangent or mirrored. These constraints reduce the degrees of freedom in the sketch. A sketch with exactly zero degrees of freedom is fully defined, while one with fewer than zero degrees of freedom is over-constrained. We generated a 2D CAD dataset derived from the SketchGraphs dataset[18], resulting in a total of 628,499 2D sketches. The statistical distribution of the number of curves and constraints operations in the 2D dataset is shown in the Figure 10.
Command Name Parameters
Constraints Series
make_horizontal(vertical) line0
fix_size curve0, size
make_coincident point0, point1
make_parallel line0, line1
make_perpendicular line0, line1
make_tangent curve0, curve1
make_mirror curve0, curve1 (same type)
make_angle line0, line1, clockwise
Table 5: CAD commands and their parameters.
Figure 10: The statistical distribution of the number of curves and constraints operations in OpenECAD 2D dataset.
@@ 6.5.2 Applications
Figure 11: An example of a real-world application of the model.
Currently, OpenECAD primarily serves as a simple CAD operation generation model, focusing on generating models from images. However, by improving the dataset, we can train models using similar methods and apply them to downstream applications (like Figure 11). These applications include, but are not limited to:
  • 1.
@@ 7 Discussion and Conclusion
In summary, we have introduced OpenECAD, a visual language model and its accompanying dataset designed for generating CAD operation sequence codes. OpenECAD aims to address the challenge of CAD model generation using visual language models and has successfully generated some relatively simple 3D shapes. We offer 5 sizes of OpenECAD models, ranging from 0.55B to 4.2B parameters.

### v3

@@ 6.6.2 Applications
Currently, OpenECAD primarily serves as a simple CAD operation generation model, focusing on generating models from images. However, by improving the dataset, we can train models using similar methods and apply them to downstream applications. These applications include, but are not limited to:
@@ 7 Discussion and Conclusion
In summary, we have introduced OpenECAD, a visual language model and its accompanying dataset designed for generating CAD operation sequence codes. OpenECAD aims to address the challenge of CAD model generation using visual language models and has successfully generated some relatively simple 3D shapes. We offer 4 sizes of OpenECAD models, ranging from 0.55B to 3.1B parameters.